# Compton Form Factor Extraction using Quantum Deep Neural Networks

### v2

@@ Abstract
We present an extraction of Compton Form Factors (CFFs) from Deeply Virtual Compton Scattering (DVCS) experiments conducted at Thomas Jefferson National Accelerator Facility, utilizing Quantum Deep Neural Networks (QDNNs). The analysis employs the standard Belitsky, Kirchner, and Müller formalism at twist-two, complemented by a fitting procedure designed to minimize model dependence in a manner analogous to conventional local fits. A pseudodata extraction test of the CFFs is performed using both Classical Deep Neural Networks (CDNNs) and QDNNs, with a detailed comparative analysis. Results indicate that QDNNs can outperform CDNNs in particular cases, offering enhanced predictive accuracy and precision even with limited model complexity. Motivated by this, we develop a metric to quantify the extent of the quantum advantage based on characteristics of DVCS experimental data. These findings underscore the promising role of QDNNs in advancing future investigations into multidimensional parton distributions and hadronic physics.
@@ I Introduction
Exploring the internal structure of the nucleon has been a longstanding goal in hadronic physics, with deeply virtual exclusive processes providing a crucial window into its partonic composition. Among these, Deeply Virtual Compton Scattering (DVCS) stands out as a uniquely sensitive probe of Generalized Parton Distributions (GPDs), which encode the spatial and momentum distributions of quarks and gluons. By linking longitudinal momentum fractions to transverse spatial coordinates, GPDs provide critical insights into the orbital angular momentum carried by partons, bridging the gap between experimental observables and the deeper theoretical understanding of nucleon structure.
Despite their fundamental importance, GPDs cannot be directly extracted from experimental measurements due to their integral appearance in DVCS amplitudes. Instead, Compton Form Factors (CFFs) emerge as a more experimentally accessible alternative, encapsulating the convolution of GPDs with perturbative coefficient functions. These complex-valued functions serve as indirect yet powerful observables, enabling the study of quark and gluon distributions through measurable quantities such as cross sections and asymmetries. This may necessitate the advancement of analysis techniques to go beyond classical methods, leveraging approaches capable of capturing the quantum correlations inherent in the DVCS process. In this context, quantum deep neural networks (QDNNs) offer a promising approach to the systematic extraction of quantum information pertaining to partonic degrees of freedom, enhancing our ability to measure and interpret the internal quantum structure of the nucleon.
QDNNs are neural network architectures that take advantage of quantum mechanics to process information. Unlike classical deep learning models that rely on traditional computation, QDNNs employ quantum circuits as trainable components, allowing them to explore complex, high-dimensional function spaces more efficiently. Libraries such as PennyLane and Quantum TensorFlow provide frameworks to construct and train QDNNs using hybrid quantum-classical approaches. These frameworks enable the encoding of classical or quantum data into quantum states, allowing manipulation through quantum circuits, and extraction of results via measurement processes. PennyLane is designed to integrate quantum computing with machine learning and automatic differentiation techniques, allowing users to build quantum nodes (or “QNodes”) that can be optimized in conjunction with classical deep learning architectures. Quantum TensorFlow extends Google’s TensorFlow to incorporate quantum computation, enabling quantum variational circuits to be trained using classical gradient descent methods. These tools provide a seamless way to incorporate quantum algorithms into machine learning tasks, facilitating the study of quantum-enhanced feature extraction, pattern recognition, and regression tasks, particularly those of a quantum nature.
One application of QDNNs is in the extraction of kinematically sensitive parameters, such as Compton Form Factors (CFFs), from helicity amplitudes in deeply virtual Compton scattering (DVCS). In DVCS, the helicity amplitudes encode critical information about the nucleon’s partonic structure, represented as complex-valued functions of kinematic variables like momentum transfer squared $t$, Bjorken scaling variable $x_{B}$, and photon virtuality $Q^{2}$. Extracting CFFs from experimental data is challenging due to limited constraints and the interference with the Bethe-Heitler process, where the final state (electron, nucleon, and photon) arises from two coherent quantum amplitudes that interfere constructively or destructively. This interference, driven by quantum superposition and phase coherence, complicates the separation of DVCS contributions. Traditionally, feed-forward deep neural networks have been used to fit experimental data and extract these parameters. However, classical models may struggle to efficiently capture the complex quantum correlations [1] and interference patterns inherent in the data, especially if the data has large experimental errors. QDNNs offer a novel approach by encoding kinematic inputs into quantum states and processing them with quantum variational circuits. While the quantum mechanical properties of QDNNs, such as entanglement in their circuit architecture, are primarily used as a flexible parameterization for regression, they may provide enhanced expressive power [1] to model the intricate structures in DVCS data compared to classical networks under particular circumstances. This potential advantage warrants further exploration to determine whether QDNNs can better represent the underlying quantum correlations in DVCS processes or simply provide an alternative complementary extraction tool.
The non-perturbative structure of helicity amplitudes in processes like DVCS often involves complex transformations between Generalized Parton Distributions (GPDs), CFFs, and measurable observables. Since QDNNs operate in a Hilbert space, they inherently support the representation of complex-valued functions, making them well-suited for approximating helicity amplitudes. Specifically, quantum feature maps in PennyLane allow one to encode cross-section data into quantum states, which can then be evolved using parameterized quantum circuits to optimize for the best representation of the underlying CFFs. A variational quantum circuit (VQC) can be trained to approximate the mapping between experimental observables, such as beam-spin asymmetries, and the CFFs that define them. This process involves quantum data encoding, where experimental data, including kinematic variables, are transformed into quantum states using amplitude or basis encoding, effectively translating kinematic dependencies into quantum states. The next step is quantum circuit evolution, in which a series of quantum gates parameterized by tunable angles is applied to transform the encoded information. These gates are chosen to represent the functional relationship between helicity amplitudes and observables. Finally, measurement and training are performed by extracting relevant quantities and using hybrid quantum-classical optimization methods, such as quantum natural gradient descent, to adjust the parameters in the quantum circuit to fit the data.
By leveraging quantum entanglement and interference, QDNNs can capture nontrivial correlations between observables that are difficult for classical models to learn efficiently. In CFF extraction, classical extraction techniques have to approximate real and imaginary parts separately, possibly missing quantum interference and correlation effects that are intrinsic to the DVCS amplitude. However, a QDNN can process this information holistically by leveraging superposition and quantum state evolution, which naturally preserve these quantum correlations [1]. Additionally, quantum-enhanced feature spaces allow for a more expressive representation of the functional dependencies of helicity amplitudes. This is particularly useful in scenarios where experimental uncertainties and systematic errors obscure direct parameter extraction. Since QDNNs operate in a higher-dimensional Hilbert space, they can encode more complex functional mappings, potentially leading to more accurate determinations of observables like beam-spin and target-spin asymmetries, which are crucial in probing the nucleon structure.
With the rapid advancement of near-term quantum processors, such as those developed by IBM, Google, and Rigetti, experimental physics stands at the threshold of utilizing QDNNs for real-world data analysis. A promising approach involves integrating quantum machine learning with Monte Carlo sampling methods to systematically propagate experimental errors through the quantum model, improving the robustness of extracted parameters like CFFs. The application of Quantum DNNs in libraries like PennyLane and Quantum TensorFlow offers a new paradigm for extracting quantum-mechanical observables in nuclear and high-energy physics. By exploiting quantum feature spaces, entanglement, and parameterized quantum circuits, these models can potentially outperform classical approaches in unraveling the helicity structure of deeply virtual processes like DVCS. As quantum hardware continues to improve, QDNNs may soon become indispensable tools for analyzing quantum field-theoretic processes, leading to deeper insights into the fundamental structure of matter. This approach may enhance the extraction fidelity of critical observables and open new avenues for studying the quantum correlations embedded within the hadronic structure.
High-statistics, finely binned data from experimental facilities such as Jefferson Lab have provided insight into the valence quark region, while complementary measurements from HERA and COMPASS have explored the role of sea quarks and gluons in deeply virtual processes. These data sets serve as the basis for generating pseudodata and set initial benchmarks for developing and validating machine learning techniques, enabling systematic comparisons between CDNNs and QDNNs.
Looking ahead, future experimental programs such as the Electron-Ion Collider (EIC) will dramatically expand the kinematic reach of DVCS measurements, particularly in the sea quark- and gluon-dominated regime. The combination of these new data with advanced QDNN-based extraction techniques offers a powerful opportunity to deepen our understanding of the nucleon’s internal structure. By leveraging quantum machine learning to disentangle the intricate relationships between spatial and momentum distributions, these methods could bring us closer to achieving a fully resolved three-dimensional mapping of partonic structure, uncovering new aspects of hadronic physics that remain inaccessible to classical approaches.
In this work, we demonstrate some very basic applications of QDNNs for information extraction and develop some metrics for comparison. Then we perform some extraction testing using a standard local fit that involves fitting a single observable at fixed kinematics, across the independent variable $\phi$, the azimuthal angle between the lepton and hadron scattering planes. Local fitting relies on the chosen framework of general approximation, such as twist truncation, leading-order contributions, and the dominance of some CFFs. At the highest approximation, the parametrization of observables, such as the cross section, in terms of CFFs is not unique as it is convention-dependent since it violates the required symmetries, which are restored by contributions that are formally suppressed by powers of $1/Q$, dubbed higher-twist corrections. In this regard, the extracted CFF depends on the choice of the observable formalism [2, 3, 4, 5, 6, 7].
While local fits offer the advantage of being potentially model-independent, extracting unique CFFs for fixed kinematics remains a formidable challenge, particularly when the number of available observables is limited, reducing the constraints in the fit. Since CFFs can vary freely within a high-dimensional parameter space, standard $\chi^{2}$ minimization methods often suffer from inconsistent local minima, leading to significant uncertainties in their extraction, even in the most rigorous analyses. As experimental datasets continue to grow in both quantity and precision, particularly with the inclusion of multiple polarized observables, the issue of data sparsity is gradually improving. However, to fully capitalize on these advances, it is equally crucial to refine and enhance extraction techniques, ensuring that they are capable of handling the increasing complexity and correlations of the data.
We employ a straightforward approach using simple feedforward architectures and, similar to previous work  [8, 9], incorporate the helicity amplitude formalism directly into the loss function during the backpropagation of error. To account for experimental uncertainties, we propagate errors through a Monte Carlo ensemble of replicas [10], where the cross section data is resampled within its experimental uncertainty. This ensemble generates a distribution of extracted CFFs, with the width of the distribution reflecting both the inherent experimental errors and the uncertainties arising from the fitting procedure.
We focus on the least constrained observable case: the unpolarized DVCS cross section. We adopt the formalism developed by Belitsky and Müller, specifically the BKM10 framework  [4], which defines the structure of the helicity amplitudes used in the extraction. While this approach is tailored to the BKM10 formalism, the extraction method itself is general and can be adapted to alternative theoretical frameworks as needed. Future work will explore an expanded approach aimed at minimizing potential biases introduced by the choice of helicity amplitude formalism. A brief overview of the BKM10 formalism will be provided in the next section.
In this study, we compare the performance of CDNNs and QDNNs in the extraction of CFFs, using a consistent theoretical framework and a new quantum metric. To ensure a direct comparison, we use the BKM10 formalism  [4] for both approaches, maintaining uniformity in the helicity amplitude structure used in the loss function. We also focus on quantifying and matching model complexity. By applying identical pseudo input data and theoretical constraints, we systematically evaluate how QDNNs enhance the extraction process relative to CDNNs, particularly in the context of the various contributions to the extraction error and fit ambiguities. The BKM10 formalism, which will be briefly discussed in the next section, provides a well-established foundation for this comparative study, allowing us to isolate the advantages or disadvantages introduced by quantum machine learning techniques.
The remainder of this article is organized as follows. Section II provides a comprehensive overview of the theoretical framework for Deeply Virtual Compton Scattering (DVCS), detailing the role of Compton Form Factors (CFFs) in the extraction process. Section III introduces Quantum Deep Neural Networks (QDNNs), explaining their architecture and advantages over Classical Deep Neural Networks (CDNNs). Examples are provided that explore quantum classification and regression, comparing QDNN and CDNN performance on basic classification and regression tasks with well-defined metrics using pseudodata. Section IV describes the experimental data from Thomas Jefferson National Accelerator Facility used both to generate pseudodata for testing and as input for the extraction of real experimentally determined CFFs in Sec. VII. Section V details the methodology for CFF extraction, including a comparative analysis of CDNN and QDNN performance using accuracy, precision, and error metrics. Section VI develops a DVCS “quantum qualifier,” a metric designed to predict whether a QDNN will outperform a CDNN in CFF extraction based purely on the characteristics of the experimental DVCS data. Section VII applies the optimized QDNN methodology to real experimental DVCS data, showcasing the practical utility and robustness of the quantum approach in a realistic setting. Finally, Sec. VIII discusses the results and concludes with key findings and future research directions for quantum-enhanced parameter extraction in hadronic physics.
@@ II Theory Framework
To fully appreciate the challenges of extracting CFFs and how quantum models might help, we begin with a review of the theoretical framework that connects GPDs, CFFs, and measurable DVCS observables. The DVCS process studies the virtual Compton scattering in deep-inelastic kinematics at the leading order in perturbative QCD. DVCS is dominated by single-quark scattering, and therefore the amplitude can be expressed in terms of the off-forward parton distributions. In the helicity-independent case, a virtual photon scatters from an unpolarized electron beam of energy $k$ off an unpolarized proton leading to the 4-fold photon electroproduction differential cross section:
$\frac{d^{4}\sigma}{dx_{B}dQ^{2}d|t|d\phi}=\frac{\alpha^{3}x_{B}y^{2}}{8\pi Q^{4}\sqrt{1+\epsilon^{2}}}\frac{1}{e^{6}}\big{|}\mathcal{T}\big{|}^{2}\;.$ (1)
The phase space of this process is parameterized by the Bjorken variable, $x_{B}=Q^{2}/2pq$, the four-momentum $q=k-k^{\prime}$ carried by the virtual photon of mass $Q^{2}={-q}^{2}=-(k-k^{\prime})^{2}$, the squared momentum transfer between the initial and final protons $t=\Delta^{2}$ with $\Delta=p^{\prime}-p$ and the lepton energy loss $y=(p\cdot q)/(p\cdot k)$. The azimuthal angle $\phi$ between the leptonic and hadronic planes is defined in the Trento convention [11]. In the cross section, $\alpha=e^{2}/(4\pi)$ is the fine structure constant, and $\epsilon=2x_{B}M/Q^{2}$ where $M$ is the proton mass.
The photon electroproduction cross section is sensitive to the coherent interference of the DVCS amplitude with the Bethe-Heitler (BH) amplitude. The DVCS process, namely, the electroproduction of a photon off a proton with the exchange of a photon with virtuality, $Q^{2}$, much larger than the momentum transfer squared $t$, hypothetically provides one of the cleanest probes of the GPDs. The BH is an easily calculated electromagnetic process, where a real photon is emitted by the initial or the final lepton. Therefore, both DVCS and BH channels contain the same final states, which can not be distinguished experimentally, and the total cross section arises from the modulus squared of the coherent sum of amplitudes:
@@
with the interference term $\mathcal{I}$ being:
@@
The BH contribution is computed exactly in [2]. It is well known from QED calculations with great precision in terms of the proton electromagnetic form factors in the range of $-t<0.4$ GeV2 and will not be discussed in detail. The unpolarized BH amplitude is given by:
T_BH^2 = e6x2By2(1 + ϵ2)2t P1(ϕ) P2(ϕ) ∑n=02 cBHn cos(nϕ).
The harmonic terms $c^{BH}_{n}$ of the BH amplitude squared depend only upon bilinear combinations of the electromagnetic form factors $F_{1}(t)$ and $F_{2}(t)$, which are computed from Kelly’s parametrization  [12]. The factors $P_{1}(\phi)$ and $P_{2}(\phi)$ are the electron propagators in the BH amplitude.
@@
At leading order (LO) in perturbative QCD and leading twist (LT) approximation, four independent proton GPDs can be accessed in the DVCS process: $H$, $E$, $\widetilde{H}$ and $\widetilde{E}$. These four GPDs depend on three variables: $x$, $\xi$ and $t$ where the latter two are experimental parameters with the skewness $\xi=x_{B}/(2-x_{B})$ at leading twist and $x\pm\xi$ the light-cone longitudinal momentum fraction of the struck quark before (“$+$” sign) and after (“--” sign) the scattering, such that the difference in the momentum fraction of the quark before absorbing the virtual photon and after the emission of the real photon is given by $2\xi$, resulting in a squared four-momentum transfer $t$ to the nucleon. This gives rise to the division of the GPDs into two distinct regions, each carrying entirely different physical interpretations. In the region where the momentum fraction $x$ exceeds $\xi$ or falls below $-\xi$ the GPDs are interpreted as the amplitudes governing the emission and reabsorption of a parton. Conversely, within the region encompassing $-\xi<x<\xi$, GPDs resemble distribution amplitudes governing the emission or absorption of a parton-antiparton pair.

### v3

@@ Abstract
We extract Compton form factors (CFFs) from deeply virtual Compton scattering measurements at the Thomas Jefferson National Accelerator Facility (JLab) using quantum‑inspired deep neural networks (QDNNs). The analysis implements the twist‑2 Belitsky–Kirchner–Müller formalism and employs a fitting strategy that emulates standard local fits. Using pseudodata, we benchmark QDNNs against classical deep neural networks (CDNNs) and find that QDNNs often deliver higher predictive accuracy and tighter uncertainties at comparable model complexity. Guided by these results, we introduce a quantitative selection metric that indicates when QDNNs or CDNNs are optimal for a given experimental fit. After obtaining local extractions from the JLab data, we perform a standard neural‑network global CFF fit and compare with previous global analyses. The results support QDNNs as an efficient and complementary tool to CDNNs for CFF determination and for future multidimensional studies of parton distributions and hadronic structure.
@@ I Introduction
Mapping the quark and gluon structure of hadrons is a central goal in QCD phenomenology. Classic determinations of parton distribution functions (PDFs) have charted the longitudinal momentum structure, but a multidimensional description that correlates longitudinal momentum with transverse spatial information remains a key objective at global accelerator facilities, and for the upcoming Electron–Ion Collider (EIC) [94, 69, 68, 99, 100]. Generalized Parton Distributions (GPDs) unify PDFs and elastic form factors in a common framework, encode correlations between transverse spatial structure and intrinsic longitudinal motion, connect to the total angular momentum via Ji’s sum rule, and provide access to mechanical properties such as pressure and shear through matrix elements of the energy–momentum tensor [4, 28, 41, 45, 97, 96].
Experimentally, GPDs are accessed in hard exclusive reactions, notably deeply virtual Compton scattering (DVCS), deeply virtual meson production (DVMP), time‑like Compton scattering (TCS), and double deeply virtual Compton scattering (DDVCS) [68, 99, 22, 59]. At leading twist the nucleon structure is encoded in four chiral‑even and four chiral‑odd quark GPDs (with analogous gluon GPDs) as functions of $(x,\xi,t)$, where $x\pm\xi$ denote the struck‑parton momentum fractions and $t$ is the momentum transfer; the physical interpretation separates into DGLAP and ERBL regions [42, 66, 43, 55, 87, 15, 46, 86]. QCD factorization theorems establish that the DVCS amplitude factorizes into a hard perturbative kernel and nonperturbative GPDs in the Bjorken limit [36, 67]. Because $x$ is not directly observable, phenomenology proceeds through Compton Form Factors (CFFs), which are convolutions of GPDs with calculable coefficient functions at leading order (LO) and beyond [19]. The real and imaginary parts of the CFFs probe principal‑value integrals and the lines $x=\pm\xi$, respectively, and their recovery from data constitutes an ill‑posed deconvolution problem [24, 90] with additional $Q^{2}$ dependence from QCD evolution [94].
Over the past two decades, a broad experimental program has mapped DVCS over complementary kinematics. H1 and ZEUS at HERA measured unpolarized cross sections and beam‑charge asymmetries at small $x_{B}$ [13, 1, 34]; HERMES provided the most complete fixed‑target set in the intermediate region [6, 7, 5, 8, 9, 10, 11]; at Jefferson Lab, CLAS and Hall A produced finely binned cross sections and polarization observables at large $x_{B}$ in the valence region, including neutron DVCS for flavor separation [105, 31, 74, 39, 40, 51, 21]; and COMPASS has reported cross sections with fixed targets at lower $x_{B}$ [12] and more recently [14] for measurement of the hard exclusive $\pi^{0}$ production cross section. New data at JLab12 (CLAS12) and future measurements at AMBER, EIC, EicC, and LHeC will greatly extend coverage and precision [49, 26, 35, 3, 77, 16, 48].
CFF extractions have followed two complementary strategies. Global fits model the xx‑dependence via GPD ansätze (e.g., VGG, GK, KM) or parameterize CFFs over all kinematics, enabling interpolation/extrapolation but inheriting model bias [84, 81, 37, 54, 106, 78]. Local fits determine CFFs pointwise in $(x_{B},t,Q^{2})$ with minimal functional assumptions, at the cost of limited predictive power [60, 58, 62, 61, 27, 85, 93, 45, 51]. To reduce modeling bias while retaining flexibility, neural‑network approaches have been developed for GPDs/CFFs that respect theory constraints and propagate experimental uncertainties through replica methods [83, 92, 37, 56, 101]; see Ref. [82] for a review. First‑principles information from lattice QCD is beginning to provide additional constraints [71, 72, 70, 63].
In this work we explore whether quantum-inspired deep neural networks (QDNNs)—simulated hybrid quantum-classical architectures—can improve CFF extractions under realistic conditions. QDNNs employ parameterized quantum circuits as trainable layers in hybrid quantum–classical models, offering expressive Hilbert‑space feature maps and entanglement that may capture interference patterns and nontrivial correlations less efficiently represented by classical networks [18]. We implement QDNNs using PennyLane’s differentiable quantum framework, enabling end‑to‑end training with classical optimizers [23]. To enable a like‑for‑like comparison, both the classical DNN (CDNN) and QDNN are trained within the same helicity‑amplitude formalism—BKM10 at twist‑2—and use identical physics‑informed losses built from the DVCS/BH decomposition [20].
We adopt the unpolarized twist-2 framework used previously Ref. [30], in which the $\phi$-dependence is driven by the BH–DVCS interference through $\Re e\,\mathcal{H}$, $\Re e\,\mathcal{E}$, and $\Re e\,\widetilde{\mathcal{H}}$, together with a $\phi$-independent DVCS contribution. We benchmark a CDNN and a QDNN on Monte Carlo pseudodata replicas, introduce a data-driven DVCS quantum qualifier to select a priori the better model class in each kinematic bin based on experimental uncertainty and the nonlinearity of the $\phi$-dependent signal, apply the qualifier-guided QDNN extraction to the JLab cross section data and compare with the KM15 global analysis [81], and finally construct a global parameterization by training a separate DNN on the combined set of best-bin CDNN/QDNN local extractions.
Overall, this study benchmarks a quantum-circuit neural-network approach for CFF extraction in the sparse, noisy regime typical of DVCS measurements. The QDNN is implemented on classical hardware via a quantum simulator, using a parameterized quantum circuit as an alternative model class—and thus an alternative algorithmic route to inference—relative to a standard CDNN. We do not require a physical quantum device; using one would introduce additional device-induced uncertainties from decoherence and hardware noise. Within a common BKM10-based forward model and training procedure, we present both separate and combined CDNN/QDNN extractions from experimental cross sections, obtaining substantially reduced uncertainties on $\Re e\,\mathcal{H}$, $\Re e\,\mathcal{E}$, and $\Re e\,\widetilde{\mathcal{H}}$ relative to representative previous extractions in the literature.
The remainder of this paper is organized as follows. In Sec. II we summarize the unpolarized BKM10 twist-2 formalism and define the forward model for the DVCS+BH cross section that enters the physics-informed loss. Section III introduces the CDNN and QDNN model classes, our quantum-circuit design choices, and the common training procedure implemented on a quantum simulator. In Sec. IV we describe the JLab datasets, kinematic selections, and the pseudodata/replica construction used for closure tests. Section V presents like-for-like benchmarks of CDNN and QDNN extractions on pseudodata, including comparisons between the baseline and fully optimized quantum architectures and a breakdown of precision, accuracy, and algorithmic contributions. In Sec. VI we construct and validate the DVCS quantum qualifier used to predict, a priori, which model class is expected to perform best in a given kinematic bin. Section VII applies the qualifier-guided strategy to experimental cross sections, produces local CFF extractions, and uses the combined best-bin results to build a global DNN parameterization for comparison with established analyses. We summarize our main findings and outline extensions to broader observable sets and increased theory fidelity in Sec. VIII.
@@ II Theory Framework
This section defines the photon electroproduction cross section and specifies the twist-2 Belitsky–Kirchner–Müller (BKM10) helicity-amplitude formalism used throughout the paper to compute the model prediction $F_{\rm pred}(\phi)$ that enters the physics-informed loss (Sec. III and Eq. (18)). In the unpolarized setting considered here, we treat the Bethe–Heitler (BH) contribution as known, and we constrain the $\phi$-dependence of the cross section through the BH–DVCS interference term at twist-2. Consistent with the simplified BKM10 setup used in Ref. [30], the only free quantities entering our forward model are $\Re e\,\mathcal{H}(x_{B},Q^{2},t)$, $\Re e\,\mathcal{E}(x_{B},Q^{2},t)$, $\Re e\,\widetilde{\mathcal{H}}(x_{B},Q^{2},t)$, and $\mathrm{DVCS}(x_{B},Q^{2},t)$, where $\mathrm{DVCS}$ parametrizes the (twist-2) $\phi$-independent DVCS-squared contribution (details below).
In the helicity-independent (unpolarized beam, unpolarized target) case, the 4-fold photon electroproduction differential cross section can be written as
$\frac{d^{4}\sigma}{dx_{B}dQ^{2}d|t|d\phi}=\frac{\alpha^{3}x_{B}y^{2}}{8\pi Q^{4}\sqrt{1+\epsilon^{2}}}\frac{1}{e^{6}}\big|\mathcal{T}\big|^{2}\;,$ (1)
with the standard DIS/DVCS variables $x_{B}=Q^{2}/(2\,p\!\cdot\!q)$, $q=k-k^{\prime}$ and $Q^{2}\equiv-q^{2}$, $t=\Delta^{2}$ with $\Delta=p^{\prime}-p$, and $y=(p\cdot q)/(p\cdot k)$. The azimuthal angle $\phi$ between the leptonic and hadronic planes is defined in the Trento convention [17]. We use $\alpha=e^{2}/(4\pi)$ and $\epsilon^{2}\equiv 4x_{B}^{2}M^{2}/Q^{2}$.
The measured final state is shared by deeply virtual Compton scattering (DVCS) and the Bethe–Heitler (BH) process. Experimentally the channels are indistinguishable, so the cross section is governed by the coherent sum of amplitudes,
@@
with interference term
@@
In our extraction, Eq. (2) is the starting point for the forward model used to compute $F_{\rm pred}$ in the loss.
The BH contribution is computed exactly in Ref. [19] and depends only on QED kinematics and the elastic proton form factors. In particular, for the unpolarized case one may write
$\big|\mathcal{T}_{BH}\big|^{2}=\displaystyle\frac{e^{6}}{x_{B}^{2}y^{2}(1+\epsilon^{2})^{2}\;t\;\mathcal{P}_{1}(\phi)\mathcal{P}_{2}(\phi)}\sum\limits_{n=0}^{2}c^{BH}_{n}\cos(n\phi)$,
where $c^{BH}_{n}$ are functions of $F_{1}(t)$ and $F_{2}(t)$ (taken from Kelly’s parametrization [76]) and $\mathcal{P}_{1,2}(\phi)$ are the lepton propagators. Practically, $|\mathcal{T}_{BH}|^{2}$ is treated as a known piece of the forward model and introduces no fit parameters.
At leading twist, QCD factorization theorems [67, 108] imply that DVCS factorizes into a perturbatively calculable hard coefficient function and nonperturbative GPDs. Because the struck-parton momentum fraction $x$ is not directly observable, DVCS observables are expressed in terms of CFFs, i.e., convolutions of GPDs with the hard kernel.
For completeness and to fix conventions, at leading order the twist-2 quark CFFs ${,,~,~}\mathcal{F}=\{\mathcal{H}$$\mathcal{E},\widetilde{\mathcal{H}}$$\widetilde{\mathcal{E}}\}$ are related to the corresponding quark GPDs $F=\{H,E,\widetilde{H},\widetilde{E}\}$ through
$\mathcal{F}(\xi,t)\equiv\sum\limits_{q}\int_{-1}^{1}dx\,C_{q}^{[\mp]}(x,\xi)\,F_{q}(x,\xi,t),$ (5)
with
$C_{q}^{[\mp]}=e_{q}^{2}\left(\frac{1}{\xi-x-i0}\mp\frac{1}{\xi+x-i0}\right),$ (6)
and $\xi=x_{B}/(2-x_{B})$ at leading twist. Each CFF is complex,
$\mathcal{F}(\xi,t)=\Re e\,\mathcal{F}(\xi,t)+i\,\Im m\,\mathcal{F}(\xi,t).$ (7)
Importantly for this work: we do not attempt a GPD deconvolution or evaluate Eq. (5) in the fit. Instead, the relevant CFF components are treated as effective, kinematic-dependent quantities to be regressed from data via the physics-informed forward model defined below.
We work in the BKM10 formalism [20] and restrict to a leading-twist regime by imposing $Q^{2}>1.5~\mathrm{GeV}^{2}$ and $|t|/Q^{2}\leq 0.25$ (consistent with a twist-2 treatment [37]). Within this setup:
  • The DVCS-squared contribution is $\phi$-independent at twist-2 for unpolarized kinematics. While the full expression can be written as a bilinear combination of CFFs, in this paper we absorb the entire DVCS-squared term into a single effective constant $\mathrm{DVCS}(x_{B},Q^{2},t)$ to be fitted.
  • The interference term provides the dominant $\phi$-dependence and, under the helicity-conserving/twist-2 restriction adopted here, depends only on $\Re e\,\mathcal{H}$, $\Re e\,\mathcal{E}$, and $\Re e\,\widetilde{\mathcal{H}}$.
Concretely, the unpolarized interference term has the harmonic structure
with Fourier coefficients
$c^{\mathcal{I}}_{n}$ $=C^{n}_{++}\,\Re e\,C^{\mathcal{I},n}_{++}(\mathcal{F})$
$\quad+C^{n}_{0+}\,\Re e\,C^{\mathcal{I},n}_{0+}(\mathcal{F}_{\mathrm{eff}})$
$\quad+C^{n}_{-+}\,\Re e\,C^{\mathcal{I},n}_{-+}(\mathcal{F}_{T})\,.$ (9)
We neglect the double helicity-flip gluonic amplitudes $C^{\mathcal{I},n}_{-+}(\mathcal{F}_{T})$ (formally suppressed by $\alpha_{s}$) and we drop helicity-changing/twist-3 effects contained in the effective CFFs $\mathcal{F}_{eff}$. With this choice, only the helicity-conserving amplitudes remain,
$\Re e\,C^{\mathcal{I},n}_{++}(\mathcal{F})$ $=\Re e\,C^{\mathcal{I}}(\mathcal{F})$
$\quad+\frac{C^{V,n}_{++}}{C^{n}_{++}}\,\Re e\,C^{\mathcal{I},V}(\mathcal{F})$
$\quad+\frac{C^{A,n}_{++}}{C^{n}_{++}}\,\Re e\,C^{\mathcal{I},A}(\mathcal{F})\,.$ (10)
where the kinematic coefficients are given in Ref. [20], and the linear CFF combinations entering the interference are
$C^{\mathcal{I}}(\mathcal{F})=F_{1}\,\mathcal{H}+\xi(F_{1}+F_{2})\,\widetilde{\mathcal{H}}-\frac{t}{4M^{2}}\,F_{2}\,\mathcal{E}$, (11)
$C^{\mathcal{I},V}(\mathcal{F})=\frac{x_{B}}{2-x_{B}+x_{B}\frac{t}{Q^{2}}}(F_{1}+F_{2})\,(\mathcal{H}+\mathcal{E}),$ (12)
$C^{\mathcal{I},A}(\mathcal{F})=\frac{x_{B}}{2-x_{B}+x_{B}\frac{t}{Q^{2}}}(F_{1}+F_{2})\,\widetilde{\mathcal{H}}$. (13)
With the restrictions above, the $\phi$-dependence of the unpolarized cross section is governed by the interference term through the three real CFFs $\Re e\,\mathcal{H}$, $\Re e\,\mathcal{E}$, and $\Re e\,\widetilde{\mathcal{H}}$, while the DVCS-squared contribution is represented by a single $\phi$-independent parameter $\mathrm{DVCS}$.
To make the connection to the remainder of the paper explicit, the model prediction entering the loss (Eq. (18)) can be viewed schematically as
$F_{\rm pred}(x_{B},Q^{2},t,\phi)$ $=F_{\rm BH}(x_{B},Q^{2},t,\phi)$
@@
$\quad+\mathrm{DVCS}(x_{B},Q^{2},t)\,.$ (14)
where $F_{\rm BH}$ is computed from the known BH expression and form factors, $F_{\rm INT}$ is computed from Eqs. (8)–(13), and $\mathrm{DVCS}$ is fitted. This is the exact simplified BKM10 formalism used in Ref. [30] and adopted here to enable a direct comparison of CDNN and QDNN extractions within the same physics-informed forward model.
III Quantum Deep Neural Networks
In this work, QDNNs are used as compact, trainable feature maps for the local regression of the three real CFFs and the $\phi$‑independent DVCS term at twist‑2,
[…]
which then enter the BKM10 helicity‑amplitude formalism of Sec. II to predict the unpolarized cross section via Eq. 2. The QDNNs are implemented with PennyLane [23] and trained end‑to‑end on a state‑vector simulator, allowing a like‑for‑like benchmark against a CDNN under the same physics‑informed loss (Sec. V). Unless stated otherwise we use $n=6$ qubits (a $2^{6}$‑dimensional embedding; cf. Fig. 2).